OpenStudy (anonymous):

You roll a blue die and a yellow die. What is P(the sum of the dice is at least 6 | the blue die shows a 3)?

OpenStudy (kropot72):

The question can be written as 'Given that the blue die shows a 3, what is the probability that the sum of the dice is 6, 7, 8 or 9' The probability of rolling a 3, 4, 5 or 6 on the yellow die is \[P(3,4,5\ or\ 6)=\frac{1}{6}+\frac{1}{6}+\frac{1}{6}+\frac{1}{6}=you\ can\ calculate\]

OpenStudy (anonymous):

Alright, thank you!